Draw and name all of the structural isomers that are ketones with five carbon atoms in its longest chain and the molecular formula C5H10O. Can this molecular formula also have an aldehyde structure? If so, illustrate and name the aldehyde. Can this be drawn as an ether? Explain.
1
Expert's answer
2012-10-16T08:58:00-0400
1)& CH3-C-CH2-CH2-CH3 2-pentanone
I I
O

2) CH3-CH2-C-CH2-CH3 & 3-pentanone
I I
O

3) CH3-C-CH-CH3 Methyl isopropyl ketone
& I I I
& O CH3


1')& CH3-CH2-CH2-CH2-HC=O& pentanal

2')& CH3-CH2-O-CH=CH-CH3